A card is drawn at a random from a pack of 52 playing cards (without jokers). Find the probabilities of the
following events happening:
A king or a red card is drawn

Answer by math_tutor2020(3816)


Answer: 7/13

----------------------------------
Work Shown:

n(A) = number of items in set A
n(Kings) = 4
n(red cards) = 26 ...... 13 hearts + 13 diamonds
n(red AND king) = 2

n(A or B) = n(A) + n(B) - n(A and B)
n(red OR king) = n(red) + n(kings) - n(red AND king)
n(red OR king) = 26 + 4 - 2
n(red OR king) = 28
28 cards are either red, a king, or both.

P(A) = probability event A happens
P(A) = n(A)/n(S) .... S is the sample space
P(red OR king) = n(red OR king)/n(any card)
P(red OR king) = 28/52
P(red OR king) = (4*7)/(4*13)
P(red OR king) = 7/13